OpenStudy (anonymous):

sec x tan x(1-sin^2x)= _____x

OpenStudy (anonymous):

write everything in terms of sine and cosine, but first start with \[1-\sin^2(x)=\cos^2(x)\]

OpenStudy (anonymous):

\[\frac{1}{\cos(x)}\times \frac{\sin(x)}{\cos(x)}\times \cos^2(x)\] and cancel away